| Entry
| Date
| Title
| Site
| Author
|
| 1 | Mon 24-Jul-1995 | Data Retrieval | 3 | Log, Paradox
|
| 2 | Thu 15-Jun-1995 | Logbook created | none | tklog
|
| 3 | Thu 15-Jun-1995 | This is the beginning log for PAMSCMS95 | none | Semmer, Steve
|
| 4 | Thu 15-Jun-1995 | Installation of PAMIII soil sensors. | none | Delany, Tony
|
| 5 | Thu 15-Jun-1995 | Dig and cook soil moisture measurements | none | Delany, Tony
|
| 6 | Thu 6-Jul-1995 | Go | 2 | Log, Paradox
|
| 7 | Sat 8-Jul-1995 | Setup | 1 | Log, Paradox
|
| 8 | Sun 09-Jul-1995 | PAM site locations | none | Horst, Tom
|
| 9 | Sun 09-Jul-1995 | ISS site access | none | Horst, Tom
|
| 10 | Sun 9-Jul-1995 | Setup | 2 | Log, Paradox
|
| 11 | Mon 10-Jul-1995 | Prop-vane orientation, PAM#2, NASA site | 2 | Horst, Tom
|
| 12 | Mon 10-Jul-1995 | Sonic anemometer, PAM#2, NASA site | 2 | Horst, Tom
|
| 13 | Mon 10-Jul-1995 | Setup | 2 | Log, Paradox
|
| 14 | Tue 11-Jul-1995 | Setup | 3 | Log, Paradox
|
| 15 | Tue 11-Jul-1995 | Setup | 1 | Log, Paradox
|
| 16 | Tue 11-Jul-1995 | PAM #3 and #1 Sonic orientation, level and calibration | none | Horst, Tom
|
| 17 | Tue 11-Jul-1995 | PAM tower locations (GPS) | none | Horst, Tom
|
| 18 | Tue 11-Jul-1995 | PAM #1 calibrations and disaster | 1 | Horst, Tom
|
| 19 | Tue 11-Jul-1995 | PAM #3 prop-vane orientation | 3 | Horst, Tom
|
| 20 | Tue 11-Jul-1995 | Soil sensors installed at PAM #1 | 1 | Michaelis, Matt
|
| 21 | Tue 11-Jul-1995 | Changed config at PAM #1. Installed "reasonable" BPH coefficients | 1 | Michaelis, Matt
|
| 22 | Tue 11-Jul-1995 | Driving Instructions to stations | none | Michaelis, Matt
|
| 23 | Wed 12-Jul-1995 | Phone numbers for site custodians | none | Horst, Tom
|
| 24 | Thu 13-Jul-1995 | wind problem | 1 | Log, Paradox
|
| 25 | Thu 13-Jul-1995 | Station dead | 3 | Log, Paradox
|
| 26 | Fri 14-Jul-1995 | sonic problem | 2 | Log, Paradox
|
| 27 | Fri 14-Jul-1995 | Sta 1 : Lightning Spike / Anem Added, RAD level | 1 | Militzer, John
|
| 28 | Fri 14-Jul-1995 | EVE_RF and EVE_FILE_TRANSLATOR working | none | Militzer, John
|
| 29 | Sat 15-Jul-1995 | Power problem | 2 | Log, Paradox
|
| 30 | Sat 15-Jul-1995 | sonic problem | 3 | Log, Paradox
|
| 31 | Sun 16-Jul-1995 | Power problem | 2 | Log, Paradox
|
| 32 | Tue 18-Jul-1995 | sonic problem | 3 | Log, Paradox
|
| 33 | Tue 18-Jul-1995 | Power problem | 2 | Log, Paradox
|
| 34 | Tue 18-Jul-1995 | wind problem | 3 | Log, Paradox
|
| 35 | Wed 19-Jul-1995 | No good reason | 1 | Log, Paradox
|
| 36 | Wed 19-Jul-1995 | Scaling for bandpass parameters | none | Horst, Tom
|
| 37 | Wed 19-Jul-1995 | Communication problem | 1 | Log, Paradox
|
| 38 | Thu 20-Jul-1995 | Communication problem | 2 | Log, Paradox
|
| 39 | Thu 20-Jul-1995 | Data Retrieval | 3 | Log, Paradox
|
| 40 | Fri 21-Jul-1995 | Dailyplots critique of data for July 20 | none | Horst, Tom
|
| 41 | Fri 21-Jul-1995 | Created netCDF data files for July 11,12,14-18 | none | Horst, Tom
|
| 42 | Sat 22-Jul-1995 | Removing despiker from EVE | none | Horst, Tom
|
| 43 | Sat 22-Jul-1995 | Moved RF antenna, fixed MRI | 1 | Horst, Tom
|
| 44 | Sat 22-Jul-1995 | sonic problem | 3 | Log, Paradox
|
| 45 | Sat 22-Jul-1995 | EVE configuration changes to date | none | Michaelis, Matt
|
| 46 | Sat 22-Jul-1995 | Setting prop parameters with the bad prop-vane software. | none | Michaelis, Matt
|
| 47 | Sat 22-Jul-1995 | Dailyplots critique for July 21 | none | Horst, Tom
|
| 48 | Sun 23-Jul-1995 | sonic blevel, PAM#1 | 1 | Horst, Tom
|
| 49 | Sun 23-Jul-1995 | PAM base location | none | Horst, Tom
|
| 50 | Sun 23-Jul-1995 | Station dead | 3 | Log, Paradox
|
| 51 | Sun 23-Jul-1995 | Brush cleared around pyranometer | 3 | Horst, Tom
|
| 52 | Sun 23-Jul-1995 | Estimate of fetch required for a/c flux measurements | none | Horst, Tom
|
| 53 | Sun 23-Jul-1995 | Data Retrieval | 3 | Log, Paradox
|
| 54 | Mon 24-Jul-1995 | Station dead | 3 | Log, Paradox
|
| 55 | Mon 24-Jul-1995 | Weckwerth photo site 1 | none | Horst, Tom
|
| 56 | Mon 24-Jul-1995 | Propvane angle, site photos, site location, hygrothermometer | 1 | Horst, Tom
|
| 57 | Mon 24-Jul-1995 | Site photos | 3 | Horst, Tom
|
| 58 | Mon 24-Jul-1995 | Data Retrieval | 3 | Log, Paradox
|
| 59 | Tue 25-Jul-1995 | Created netCDF data files for July 19-22, 24 for station 3 | none | Horst, Tom
|
| 60 | Tue 25-Jul-1995 | Lightning hit on 7/22/95 knocked out the station | 3 | Michaelis, Matt
|
| 61 | Wed 26-Jul-1995 | Data Retrieval | 1 | Log, Paradox
|
| 62 | Wed 26-Jul-1995 | Merged direct rf ppf files into netCDF files for July 18-26 | none | Horst, Tom
|
| 63 | Wed 26-Jul-1995 | Photo recon of PAM sites | none | Horst, Tom
|
| 64 | Thu 27-Jul-1995 | sonic problem | 3 | Log, Paradox
|
| 65 | Fri 28-Jul-1995 | Data Retrieval | 2 | Log, Paradox
|
| 66 | Fri 28-Jul-1995 | sonic problem | 1 | Log, Paradox
|
| 67 | Fri 28-Jul-1995 | Merged local data storage from station 3 into neCDF files for July 24-26 | none | Horst, Tom
|
| 68 | Sat 29-Jul-1995 | Local data storage files: Status of Offload from station, translation to ppf and merging | none | Michaelis, Matt
|
| 69 | Sat 29-Jul-1995 | Instructions for ppf'ing local data storage files | none | Michaelis, Matt
|
| 70 | Sat 29-Jul-1995 | Created netCDF data files for July 19, 21-27 | none | Horst, Tom
|
| 71 | Sun 30-Jul-1995 | Installed hygrothermometer and GOES xmtr | 3 | Militzer, John
|
| 72 | Sun 30-Jul-1995 | Aux sensor | 2 | Log, Paradox
|
| 73 | Sun 30-Jul-1995 | Station in water: Logger drenched / removed | 2 | Militzer, John
|
| 74 | Mon 31-Jul-1995 | Station Electronics Taken Down - Hurricane coming! | 2 | Militzer, John
|
| 76 | Tue 01-Aug-1995 | Translation from pam ppf files to netCDF files | none | Horst, Tom
|
| 77 | Tue 1-Aug-1995 | sonic problem | 1 | Log, Paradox
|
| 78 | Thu 3-Aug-1995 | Communication problem | 3 | Log, Paradox
|
| 79 | Thu 03-Aug-1995 | Turned sonic back on | 1 | Horst, Tom
|
| 80 | Thu 3-Aug-1995 | Data Retrieval | 2 | Log, Paradox
|
| 81 | Thu 3-Aug-1995 | Station dead | 2 | Log, Paradox
|
| 82 | Fri 04-Aug-1995 | Reboot parrot | none | Semmer, Steve
|
| 83 | Fri 4-Aug-1995 | Data Retrieval | 1 | Log, Paradox
|
| 84 | Fri 04-Aug-1995 | Site visit on Thursday to PAM2 | 2 | Semmer, Steve
|
| 85 | Fri 04-Aug-1995 | Site visit on Thursday to PAM3 | 3 | Semmer, Steve
|
| 86 | Fri 04-Aug-1995 | Visit to PAM 1 | 1 | Semmer, Steve
|
| 87 | Fri 4-Aug-1995 | Power problem | 3 | Log, Paradox
|
| 88 | Fri 04-Aug-1995 | Site Visit to PAM3 | 3 | Militzer, John
|
| 89 | Sat 5-Aug-1995 | T-RH problem | 2 | Log, Paradox
|
| 90 | Sat 05-Aug-1995 | Visit to PAM 2 | 2 | Semmer, Steve
|
| 91 | Sun 06-Aug-1995 | Visit station 3 | 3 | Semmer, Steve
|
| 92 | Mon 07-Aug-1995 | Daily Status, 219 (8/7/95) | none | Semmer, Steve
|
| 93 | Mon 07-Aug-1995 | Sonic levels on PAM 2 back | none | Semmer, Steve
|
| 94 | Mon 7-Aug-1995 | Routine | 1 | Log, Paradox
|
| 95 | Mon 07-Aug-1995 | Routine Visit to Station 1 | 1 | Militzer, John
|
| 96 | Mon 7-Aug-1995 | Data Retrieval | 3 | Log, Paradox
|
| 97 | Tue 08-Aug-1995 | Visit to station 3 yesterday | 3 | Semmer, Steve
|
| 98 | Tue 8-Aug-1995 | rain gauge work | 3 | Log, Paradox
|
| 99 | Tue 8-Aug-1995 | Routine | 2 | Log, Paradox
|
| 100 | Tue 08-Aug-1995 | Visit to station 3, 9:00 to 11:00 local | 3 | Semmer, Steve
|
| 101 | Tue 08-Aug-1995 | Station 2 visit, 1:00 to 3:30 pm | 2 | Semmer, Steve
|
| 102 | Wed 09-Aug-1995 | Daily Status, 220 (8/8/95) | none | Semmer, Steve
|
| 103 | Wed 9-Aug-1995 | rad problem | 2 | Log, Paradox
|
| 105 | Wed 9-Aug-1995 | Replaced ETI | 3 | Log, Paradox
|
| 107 | Thu 10-Aug-1995 | Fetch for a/c flights | none | Horst, Tom
|
| 108 | Thu 10-Aug-1995 | Station 2 visit yesterday | 2 | Semmer, Steve
|
| 109 | Thu 10-Aug-1995 | Daily Status, 221 (8/9/95) | none | Semmer, Steve
|
| 110 | Thu 10-Aug-1995 | Data Retrieval | 1 | Log, Paradox
|
| 111 | Thu 10-Aug-1995 | tklog entries inserted for COMPAQ LOGBOOK | none | Militzer, John
|
| 112 | Thu 10-Aug-1995 | Routine | 3 | Log, Paradox
|
| 113 | Thu 10-Aug-1995 | Station 1 visit this morning | 1 | Semmer, Steve
|
| 114 | Thu 10-Aug-1995 | Station 3 visit, sonic and download | 3 | Semmer, Steve
|
| 115 | Fri 11-Aug-1995 | Daily Status, 223 (8/11/95) | none | Semmer, Steve
|
| 116 | Fri 11-Aug-1995 | sonic problem | 2 | Log, Paradox
|
| 117 | Fri 11-Aug-1995 | Station 2 visit, download and v component | 2 | Semmer, Steve
|
| 118 | Fri 11-Aug-1995 | GOES going down at station 3 | 3 | Semmer, Steve
|
| 119 | Sat 12-Aug-1995 | Daily Status, 224 (8/12/95) | none | Semmer, Steve
|
| 120 | Sat 12-Aug-1995 | Goes to be disconnected at station 2 | 2 | Semmer, Steve
|
| 121 | Sat 12-Aug-1995 | Communication problem | 2 | Log, Paradox
|
| 122 | Sat 12-Aug-1995 | New TKLOG Import function : Compaq Logbook entries added | none | Militzer, John
|
| 123 | Sun 13-Aug-1995 | Daily Status, 225 (8/13/95) | none | Semmer, Steve
|
| 124 | Sun 13-Aug-1995 | GOES noise test on sonics | none | Semmer, Steve
|
| 125 | Sun 13-Aug-1995 | Routine | 1 | Log, Paradox
|
| 126 | Sun 13-Aug-1995 | Routine | 2 | Log, Paradox
|
| 127 | Sun 13-Aug-1995 | Routine | 3 | Log, Paradox
|
| 128 | Sun 13-Aug-1995 | Site visits today | none | Semmer, Steve
|
| 129 | Sun 13-Aug-1995 | Moved BPH sensors to height of hygros at all sites | none | Semmer, Steve
|
| 130 | Sun 13-Aug-1995 | Barometer IC's Report | none | Militzer, John
|
| 131 | Sun 13-Aug-1995 | Assman IC's Report | none | Militzer, John
|
| 132 | Sun 13-Aug-1995 | Station Visitation Summary from Compaq/Paradox | none | Militzer, John
|
| 134 | Tue 15-Aug-1995 | Data Retrieval | 3 | Log, Paradox
|
| 135 | Tue 15-Aug-1995 | Data Retrieval | 1 | Log, Paradox
|
| 136 | Tue 15-Aug-1995 | Data Retrieval | 2 | Log, Paradox
|
| 137 | Tue 15-Aug-1995 | Data Retrieval | 2 | Log, Paradox
|
| 138 | Tue 15-Aug-1995 | FInal visits to sites | none | Semmer, Steve
|
| 139 | Fri 01-Sep-1995 | Backups to Stout Completed | none | Militzer, John
|
| 140 | Mon 09-Oct-1995 | Post SCMS95 pressure sensor report | | Semmer, Steve
|
| 141 | Mon 09-Oct-1995 | Post SCMS95 temperature/humidity sensor report | | Semmer, Steve
|
| 142 | Fri 27-Oct-1995 | ppf data files converted to netCDF | none | Horst, Tom
|